# Ultrafast All-Optical Measurement of Squeezed Vacuum in a Lithium Niobate Nanophotonic Circuit

### v3

@@ I Introduction
Many quantum systems have been used to gain an advantage over otherwise purely classical means in a variety of fields [1, 2, 3]. Photonics has emerged as a front-runner platform in quantum information processing (QIP) for several key reasons. Most photonic technologies are capable of operating at room temperature outside a carefully-controlled cryogenic environment. Advances in integrated photonics have allowed many devices and circuits to be combined into a single monolithic platform similar to CMOS technology and the advent of integrated circuits [4]. Photonics also offers an inherently broad bandwidth which, when combined with dispersion engineering, can allow for the manipulation and propagation of ultra-short pulses of light [5]. Time-multiplexing, a pulse-based encoding technique used in photonic systems, can leverage femtosecond pulses to scale information density and throughput, and can exceed clock speeds beyond what is currently possible with conventional electronics [6, 7, 8, 9].
@@
A crucial part of any QIP system is a measurement device capable of characterizing a component of interest. Homodyne detection is widely employed as its phase-sensitive nature can be used to reconstruct the Wigner function, a quasi-probability distribution over two non-commuting variables which completely characterizes a quantum state [10]. Homodyne detectors can also isolate a single mode of a state of interest by shaping the spectral profile of the local oscillator appropriately [11]. This phase-sensitive and mode-selective behavior makes them well-suited for characterizing a variety of quantum states, including non-classical states. Single-photon and photon number-resolving (PNR) detectors are also common tools for state characterization and have been used to generate and measure of states with Wigner negativity [12, 13, 14, 15, 16, 17]. While homodyne and PNR detection are powerful techniques for probing quantum states of interest, their speed is ultimately constrained by the bandwidth of the electronics used to physically realize these detectors, which are limited to the GHz range.
@@
To overcome the speed limitations of electronics, we demonstrate chip-scale all-optical Wigner tomography of squeezed vacuum (SV) using a nanophotonic optical parametric amplifier (OPA). In our previous work [18], we measured the average photon number at the output of an OPA to calculate squeezing levels. In this work, we use a fast photodetector to measure each pulse and resolve the statistical information necessary to recover the Wigner function of the input SV. We design our OPA to have low dispersion at both the pump and signal wavelengths, allowing for the distortion-free propagation and amplification of femtosecond pulses. Such an OPA-based circuit supports a maximum measurement repetition rate (i.e. clock speed) of 6.5 THz. Increased clock speeds offers crucial benefits for QIP in photonic time-multiplexed systems as higher clock speeds enables greater information density and shorter processing times [19, 20]. Our results highlight nanophotonic OPAs as an important building block for chip-scale ultrafast quantum information processing systems at room temperature.
@@ II Parametric Amplifiers as Quantum Measurement Devices
Early proposals of OPAs as quantum measurement devices demonstrated loss-tolerance and detector inefficiency mitigation [21]. OPAs are particularly well-suited for this task as their phase-sensitive amplification is in principle noiseless unlike phase-insensitive amplifiers such as erbium-doped fibers and semiconductor gain media [22]. OPA-assisted balanced homodyne detection has been used to characterized fiber-based sources of SV over 43 GHz of electronic bandwidth [23]. A similar technique has also been demonstrated using the $\chi^{(3)}$ nonlinearity [24]. Other experiments have shown quantum-enhanced sensing using an SU(1,1) interferometer constructed from two OPAs for state readout [25, 26]. OPAs have also been used to amplify weak spatially-varying signals and detect quantum correlations for imaging applications [27, 28]. All-optical feed-forward, a technique used to surpass the constraints of electronics in feed-forward schemes for quantum information processing, has been demonstrated using 3 OPAs [29]. Wigner-tomography of squeezed states using bulk OPAs has been experimentally realized [30]. OPA-based techniques have also been extended to squeezing across multiple spatial modes where direct detection is used to disentangle and analyze each mode independently [31].
@@ II.1 Quadrature Measurement
Figure 1 depicts the scheme used to measure the Wigner function of a SV state encoded in a femtosecond pulse. When the SV enters the nanophotonic amplifier, its in-phase quadrature $\hat{x}_{\phi}$ is amplified by the pump while its out-of-phase quadrature $\hat{p}_{\phi}$ is deamplified. These quadratures map to the original operators of the state as [30]:
@@
For a large gain, the 2nd and 3rd terms of Eq.5 can be ignored such that:
@@
Because our OPA exhibits low dispersion over a broad bandwidth while operating in the type-0 phase matching configuration, it amplifies many orthogonal spectro-temporal modes simultaneously [32], all of which contribute to the measured $P(N,\phi)$. These modes can be calculated analytically by computing the joint spectral intensity (JSI) function of the signal and idler. Our JSI and the corresponding modes are computed and plotted in the supplementary using the dispersion parameters calculated from the waveguide geometry and the Heisenberg propagators derived in [33]. Because our JSI is inseparable ($JSI(\omega_{s},\omega_{i})\neq\phi(\omega_{s})\phi(\omega_{i})$), it is composed of multiple independent modes $\Phi(\omega_{s},\omega_{i})$ of the form $\Phi(\omega_{s},\omega_{i})=\phi(\omega_{s})\phi(\omega_{i})$.
At the photodetector, we use a combination of a 1700-nm long-pass filter and 1950-nm short-pass filter to suppress contributions from higher order modes. Using the measured $P(N)$ for vacuum amplification (i.e. when no signal field is sent into the OPA), we calculate a Schmidt number of 1.35 modes after filtering from the definition of $g^{(2)}$ in [34]. Contributions from the higher-order modes are quantified and separated by fitting a 2-mode photon number distribution to determine the gain of the fundamental mode as a function of $\phi$. As our Schmidt number of 1.35 is close to single-mode (i.e. 1), we limit our fit to just the first two modes as contributions from higher order modes are relatively small. For a single mode, the photon number distribution is
@@
where $\langle N(\phi)\rangle$ is the squeezing-dependent average photon number. Convolving this distribution with itself yields the 2-mode distribution [35]:
[…]
@@ II.2 Dispersion Engineering
A key advantage of using nanophotonics is the ability to control the dispersive properties of the waveguides used to realize photonic circuits. Figure 3 depicts our ridge waveguide geometry and material stack used for our OPAs. We fabricate our OPAs on a thin-film lithium niobate (TFLN) on silica on silicon wafers available from NanoLN. By adjusting the width, etch depth, and thin-film thickness of the waveguide, we can manipulate its dispersive properties to achieve low group velocity dispersion (GVM) at the pump and signal wavelengths while simultaneously minimizing the walk-off from group velocity mismatch (GVM) between the pump and signal [5]. Figures 2A-B plot the GVD for the signal/pump and GVM at degeneracy for the waveguide geometry in Fig. 3A. The operating wavelengths are marked with a red star. This regime of operation allows for our OPA to exhibit high gain over a broad bandwidth, making it an ideal tool for studying quantum states encoded in ultrafast pulses. Our waveguide geometry achieves a GVD of $-17.3~{}\mathrm{fs^{2}/mm}$ at our signal wavelength of 1860 nm, $\mathrm{244~{}fs^{2}/mm}$ at our pump wavelength of 930 nm, and a GVM of $\mathrm{-87~{}fs/mm}$. By comparison, bulk lithium niobate has a GVD of $\mathrm{13.3~{}fs^{2}/mm}$ at 1860 nm, a GVD of $\mathrm{341~{}fs^{2}/mm}$ at 930 nm, and a GVM of $\mathrm{203~{}fs/mm}$. A thin layer of poly-methyl methacrylate (PMMA) is deposited via spin-coating on top of the waveguide which acts to perturb the effective refractive index of the waveguide and offer some tunability of the phase matching condition during fabrication. We note that our design is not perfectly optimal, and that geometries with lower dispersion have been demonstrated [5].
@@
Figure 2C shows the parametric generation (i.e. vacuum amplification) spectra for the 2.5-mm OPA used for SV generation and the 5-mm OPA used for measurement. These spectra are taken at a pump pulse energy of 30 pJ. Both OPAs exhibit a 3-dB gain over $>$ 20 THz as a result of low GVD at the signal wavelength and signal gains $>$ 40 dB thanks to low GVM. Compared to the squeezer, the measurement OPA has a slightly narrower and more intense gain spectrum as the extra 2.5-mm of propagation allows for greater walk-off far from degeneracy and more interaction time near degeneracy.
@@
Operating with low dispersion allows us to utilize ultrafast femtosecond pulses at both the pump and signal frequencies. Shorter pulses are advantageous for nonlinear processes such as OPA because the high peak power from temporal confinement can achieve a stronger nonlinear interaction for the same pulse energy. This interaction is also enhanced by the tight spatial mode confinement offered by ridge waveguides. Furthermore, by shortening the length of the pump and signal pulses in the time domain, subsequent pulses can be packed closer together in time, enabling a significant boost in clock speeds for time-multiplexing. This boost also allows larger quantum circuits to be realized in cluster state architectures [19, 20]. Based on the dispersive properties of our 5-mm measurement OPA and our 70-fs pump source, we estimate an upper bound on the temporal length of our SV of 154.3 fs, which gives a theoretical maximum of 6.5 THz for our clock rate [38]. This clock rate can be increased further by using a shorter OPA and a higher pump power, or more advanced dispersion engineering.
@@ II.3 Pump Depletion
An important consideration for OPAs as quantum measurement devices is the depletion experienced by the pump during amplification. As the pump energy increases, the nonlinear interaction becomes more efficient, causing the pump pulse to deplete while propagating in the OPA and the gain to saturate. Figure 3C shows measured on-chip OPA gain vs the input pump energy. The black line represents the theoretical undepleted pump gain $G_{dB}=10\log_{10}(e^{b\sqrt{E}})$ where $E$ is the pump energy in Joules and $b$ is a found by fitting the pump energy vs gain curve in the undepleted regime. The dashed blue line represents the theoretical gain when accounting for pump depletion up to 2nd order in the Gaussian limit [37]. Around 20 pJ, the measured gain begins to saturate relative to the theoretical gain due to pump depletion effects.
Pump depletion impacts our measurement in two ways. First, the larger photon number components of the state being measured experience less gain than smaller photon number components. This phenomenon, often known as gain saturation for classical amplifiers, suppresses the larger photon number contributions to the measured distributions and distorts our measurement. This distortion can be mitigated by improving output coupling losses to allow operation at a lower pump power. We investigate this distortion experimentally in section IV.2. Second, pump depletion can be thought of as a semi-deterministic transfer of energy from the pump pulse to the signal pulse. This causes the peak of the photon number distribution to shift away from $N=0$. We observe this effect in classical simulations which use a split-step Fourier method to simulate single-mode and multimode vacuum measurements. These simulations are detailed in the supplementary. This effect has also been observed in single-mode quantum simulations [37] and experimentally [34]. We address this issue when fitting our data with 2-mode distributions by fitting a constant offset along the $N$ axis, leading to good agreement between the measured data and fitted distributions.
@@ III Experimental Setup
Figure 4: Experimental setup. MLL: Titanium:sapphire tunable mode-locked laser. PBS: polarizing beamsplitter. VND: variable neutral density filter. SH: mechanical shutter. Obj: reflective objective. FC: reflective fiber collimator. LP and SP: long-pass and short-pass wavelength filters. 90/10: fiber splitter with 90% going to the fast detector and 10% going to the slow oscilloscope. PD: photodetector. Fast Osc: 80 GSPS 40 GHz oscilloscope. Slow Osc: 100 MSPS 10 MHz oscilloscope.
Figure 4 shows the experimental setup. An 80 MHz mode-locked titanium-sapphire laser (MLL) sends 70-fs pulses to a half-wave plate (HWP) and polarizing beamsplitter (PBS) combination to control the power splitting between the measurement and squeezer beam paths. Pump light for the squeezer OPA is sent to a delay stage used to scan $\phi$ in equation 1. A shutter is also placed in this path to block the squeezer beam for shot-noise calibration. Two continuously-variable neutral density filters (VND) are used in both paths to fine-tune the input power. The paths are combined on a reflective objective (Obj) and focused onto the waveguide inputs for the OPAs. Figure 4A shows the layout of the nanophotonic circuit. In the squeezer OPA, the vacuum field around 1860 nm is squeezed by the pump to generate squeezed vacuum (SV). The SV and the remaining pump light enter an adiabatic coupler which passes at least 55% of the SV and 5% of the pump into the measurement OPA. Inside the measurement OPA, the SV is amplified by a strong pump pulse (40 pJ) to macroscopic photon levels detectable by a fast photodetector. A lensed fiber collects light from the output of the measurement OPA and directs it to a series of free-space filters used to remove remaining pump light, including the previously mentioned 1700-nm long pass and a 1950-nm short pass for limiting the measured mode number. From here, the filtered SV is collected into single-mode fiber and enters a 90/10 splitter where 90% is sent to a fast (10 GHz) photodetector and fast (80 GSPS 40 GHz) oscilloscope used to resolve individual pulses while the remaining 10% is sent to a slow (1 MHz) detector and slow (100 MSPS 10 MHz) oscilloscope used to detect the amplification fringe and determine the measurement phase $\phi$.

### v4

@@ I Introduction
Many quantum systems have been used to gain an advantage over otherwise purely classical means in a variety of fields [1, 2, 3]. Photonics has emerged as a front-runner platform in quantum information processing (QIP) for several key reasons. Most photonic technologies are capable of operating at room temperature outside a carefully-controlled cryogenic environment. Advances in integrated photonics have allowed many devices and circuits to be combined into a single monolithic platform similar to CMOS technology and the advent of integrated circuits [4]. Photonics also offers an inherently broad bandwidth which, when combined with dispersion engineering, can allow for the manipulation and propagation of ultra-short pulses of light [5]. Time-multiplexing, a pulse-based encoding technique used in photonic systems, can leverage femtosecond pulses to scale information density and throughput, and can exceed clock speeds beyond what is currently possible with conventional electronics [6, 7, 8].
@@
A crucial part of any QIP system is a measurement device capable of characterizing a component of interest. Homodyne detection is widely employed as its phase-sensitive nature can be used to reconstruct the Wigner function, a quasi-probability distribution over two non-commuting variables which completely characterizes a quantum state [9]. Homodyne detectors can also isolate a single mode of a state of interest by shaping the spectral profile of the local oscillator appropriately [10]. This phase-sensitive and mode-selective behavior makes them well-suited for characterizing a variety of quantum states, including non-classical states. Single-photon and photon number-resolving (PNR) detectors are also common tools for state characterization and have been used to generate and measure states with Wigner negativity [11, 12, 13, 14, 15, 16]. While homodyne and PNR detection are powerful techniques for probing quantum states of interest, their speed is ultimately constrained by the bandwidth of the electronics used to physically realize these detectors, which are limited to the GHz range.
@@
To overcome the speed limitations of electronics, we demonstrate chip-scale all-optical Wigner tomography of squeezed vacuum (SV) using a nanophotonic optical parametric amplifier (OPA). In our previous work [17], we measured the average photon number at the output of an OPA to calculate squeezing levels. In this work, we use a fast photodetector to measure each pulse and resolve the statistical information necessary to recover the Wigner function of the input SV. We design our OPA to have low dispersion at both the pump and signal wavelengths, allowing for the distortion-free propagation and amplification of femtosecond pulses. Such an OPA-based circuit supports a maximum measurement repetition rate (i.e. clock speed) of 6.5 THz. Increased clock speeds offers crucial benefits for QIP in photonic time-multiplexed systems as faster clocks lead to shorter processing times while also increasing the maximum circuit size that can be realized for a given system [18, 19, 20]. Our results highlight nanophotonic OPAs as an important building block for chip-scale ultrafast quantum information processing systems at room temperature.
@@ II Parametric Amplifiers as Quantum Measurement Devices
Early proposals of OPAs as quantum measurement devices demonstrated loss-tolerance and detector inefficiency mitigation [21]. OPAs are particularly well-suited for this task as their phase-sensitive amplification is in principle noiseless unlike phase-insensitive amplifiers such as erbium-doped fibers and semiconductor gain media [22]. OPA-assisted balanced homodyne detection has been used to characterize fiber-based sources of SV over 43 GHz of electronic bandwidth [23]. A similar technique has also been demonstrated using the $\chi^{(3)}$ nonlinearity [24]. Other experiments have shown quantum-enhanced sensing using an SU(1,1) interferometer constructed from two OPAs for state readout [25, 26]. OPAs have also been used to amplify weak spatially-varying signals and detect quantum correlations for imaging applications [27, 28]. All-optical feed-forward, a technique used to surpass the constraints of electronics in feed-forward schemes for quantum information processing, has been demonstrated using three OPAs [29]. Wigner-tomography of squeezed states using bulk OPAs has been experimentally realized [30]. OPA-based techniques have also been extended to squeezing across multiple spatial modes where direct detection is used to disentangle and analyze each mode independently [31].
@@ II.1 Quadrature Measurement
Figure 1 depicts the scheme used to measure the Wigner function of an SV state encoded in a femtosecond pulse. When the SV enters the nanophotonic amplifier, its in-phase quadrature $\hat{x}_{\phi}$ is amplified by the pump while its out-of-phase quadrature $\hat{p}_{\phi}$ is deamplified. These quadratures map to the original operators of the state as [30]:
@@
For a large gain, the second and third terms of Eq.5 can be ignored such that:
@@
Because our OPA exhibits low dispersion over a broad bandwidth while operating in the type-0 phase matching configuration, it amplifies many orthogonal spectro-temporal modes simultaneously [32], all of which contribute to the measured $P(N,\phi)$. These modes can be calculated analytically by computing the joint spectral intensity (JSI) function of the signal and idler. Our JSI and the corresponding modes are computed and plotted in Appendices using the dispersion parameters calculated from the waveguide geometry and the Heisenberg propagators derived in [33]. Because our JSI is inseparable ($JSI(\omega_{s},\omega_{i})\neq\phi(\omega_{s})\phi(\omega_{i})$), it is composed of multiple independent modes $\Phi(\omega_{s},\omega_{i})$ of the form $\Phi(\omega_{s},\omega_{i})=\phi(\omega_{s})\phi(\omega_{i})$.
At the photodetector, we use a combination of a 1700-nm long-pass filter and a 1950-nm short-pass filter to suppress contributions from higher-order modes. Using the measured $P(N)$ for vacuum amplification (i.e. when no signal field is sent into the OPA), we calculate a Schmidt number of 1.35 modes after filtering from the definition of $g^{(2)}$ in Ref. [34]. Contributions from the higher-order modes are quantified and separated by fitting a two-mode photon number distribution to determine the gain of the fundamental mode as a function of $\phi$. As our Schmidt number of 1.35 is close to single-mode (i.e. 1), we limit our fit to the first two modes as contributions from higher-order modes are relatively small. For a single mode, the photon number distribution is
@@
where $\langle N(\phi)\rangle$ is the squeezing-dependent average photon number. Convolving this distribution with itself yields the two-mode distribution [35]:
$P^{[2]}(N,\phi)=\int_{0}^{\infty}P^{[1]}_{\langle N_{1}\rangle}(N-n,\phi)P^{[1]}_{\langle N_{2}\rangle}(n,\phi)dn$ (8)
@@ II.2 Dispersion Engineering
A key advantage of using nanophotonics is the ability to control the dispersive properties of the waveguides used to realize photonic circuits. Figure 3 depicts our ridge waveguide geometry and material stack used for our OPAs. We fabricate our OPAs on a thin-film lithium niobate (TFLN) on silica on silicon wafers available from NanoLN. By adjusting the width, etch depth, and thin-film thickness of the waveguide, we can manipulate its dispersive properties to achieve low group velocity dispersion (GVD) at the pump and signal wavelengths while simultaneously minimizing the walk-off from group velocity mismatch (GVM) between the pump and signal [5]. Figures 2A-B plot the GVD for the signal/pump and GVM at degeneracy for the waveguide geometry in Fig. 3A. The operating wavelengths are marked with a red star. This regime of operation allows for our OPA to exhibit high gain over a broad bandwidth, making it an ideal tool for studying quantum states encoded in ultrafast pulses. Our waveguide geometry achieves a GVD of $17.3~\mathrm{fs^{2}/mm}$ at our signal wavelength of 1860 nm, $\mathrm{244~fs^{2}/mm}$ at our pump wavelength of 930 nm, and a GVM of $\mathrm{-87~fs/mm}$. By comparison, bulk lithium niobate has a GVD of $\mathrm{13.3~fs^{2}/mm}$ at 1860 nm, a GVD of $\mathrm{341~fs^{2}/mm}$ at 930 nm, and a GVM of $\mathrm{203~fs/mm}$. A thin layer of poly-methyl methacrylate (PMMA) is deposited via spin-coating on top of the waveguide that acts to perturb the effective refractive index of the waveguide and offer some tunability of the phase matching condition during fabrication. We note that our design is not perfectly optimal, and that geometries with lower dispersion have been demonstrated [5]. Ideally, the dispersion of the squeezer and measurement OPAs should be chosen such that the overlap of their fundamental modes is maximized while also minimizing the GVM and GVD. These goals are not always mutually compatible, and hence there is some trade-off. For this work, we focus on minimizing GVM and GVD while relying on filtering to enforce mode overlap.
@@
Figure 2C shows the parametric generation (i.e. vacuum amplification) spectra for the 2.5-mm OPA used for SV generation and the 5-mm OPA used for measurement. These spectra are taken at a pump pulse energy of 30 pJ. Both OPAs exhibit a 3-dB gain over >> 20 THz as a result of low GVD at the signal wavelength and signal gains >> 40 dB thanks to low GVM. Compared to the squeezer, the measurement OPA has a slightly narrower and more intense gain spectrum as the extra 2.5-mm of propagation allows for greater walk-off far from degeneracy and more interaction time near degeneracy.
@@
Operating with low dispersion allows us to utilize ultrafast femtosecond pulses at both the pump and signal frequencies. Shorter pulses are advantageous for nonlinear processes such as OPA because the high peak power from temporal confinement can achieve a stronger nonlinear interaction for the same pulse energy. This interaction is also enhanced by the tight spatial mode confinement offered by ridge waveguides. Furthermore, by shortening the length of the pump and signal pulses in the time domain, subsequent pulses can be packed closer together in time, enabling a significant boost in clock speeds for time-multiplexing. This boost also allows larger quantum circuits to be realized in cluster state architectures [18, 19, 20]. Based on the dispersive properties of our 5-mm measurement OPA and our 70-fs pump source, we estimate an upper bound on the temporal length of our SV of 154.3 fs, which gives a theoretical maximum of 6.5 THz for our clock rate [38]. This clock rate can be increased further by using a shorter OPA and a higher pump power, or more advanced dispersion engineering.
@@ II.3 Pump Depletion
An important consideration for OPAs as quantum measurement devices is the depletion experienced by the pump during amplification. As the pump energy increases, the nonlinear interaction becomes more efficient, causing the pump pulse to deplete while propagating in the OPA and the gain to saturate. Figure 3C shows the measured on-chip OPA gain vs the input pump energy. The black line represents the theoretical undepleted pump gain $G_{dB}=10\log_{10}(e^{b\sqrt{E}})$ where $E$ is the pump energy in Joules and $b$ is a found by fitting the pump energy vs gain curve in the undepleted regime. The dashed blue line represents the theoretical gain when accounting for pump depletion up to second order in the Gaussian limit [37]. Around 20 pJ, the measured gain begins to saturate relative to the theoretical gain due to pump depletion effects.
Pump depletion impacts our measurement in two ways. First, the larger photon number components of the state being measured experience less gain than smaller photon number components. This phenomenon, often known as gain saturation for classical amplifiers, suppresses the larger photon number contributions to the measured distributions and distorts our measurement. This distortion can be mitigated by improving output coupling losses to allow operation at a lower pump power. We investigate this distortion experimentally in section IV.2. Second, pump depletion can be thought of as a semi-deterministic transfer of energy from the pump pulse to the signal pulse. This causes the peak of the photon number distribution to shift away from $N=0$. We observe this effect in classical simulations that use a split-step Fourier method to simulate single-mode and multimode vacuum measurements. These simulations are detailed in the supplementary. This effect has also been observed in single-mode quantum simulations [37] and experimentally [34]. We address this issue when fitting our data with 2-mode distributions by fitting a constant offset along the $N$ axis, leading to good agreement between the measured data and fitted distributions.
@@ III Experimental Setup
Figure 4: Experimental setup. MLL: Titanium:sapphire tunable mode-locked laser. PBS: polarizing beam splitter. VND: variable neutral density filter. SH: mechanical shutter. Obj: reflective objective. FC: reflective fiber collimator. LP and SP: long-pass and short-pass wavelength filters. 90/10: fiber splitter with 90% going to the fast detector and 10% going to the slow detector. PD: photodetector. Fast Osc: 80 GSPS 40 GHz oscilloscope. Slow Osc: 100 MSPS 10 MHz oscilloscope.
Figure 4 shows the experimental setup. An 80 MHz mode-locked titanium-sapphire laser (MLL) sends 70-fs pulses to a half-wave plate (HWP) and polarizing beam splitter (PBS) combination to control the power splitting between the measurement and squeezer beam paths. Pump light for the squeezer OPA is sent to a delay stage used to scan $\phi$ in equation 1. A shutter is also placed in this path to block the squeezer beam for shot-noise calibration. Two continuously-variable neutral density filters (VNDs) are used in both paths to fine-tune the input power. The paths are combined on a reflective objective (Obj) and focused onto the waveguide inputs for the OPAs. Figure 4A shows the layout of the nanophotonic circuit. In the squeezer OPA, the vacuum field around 1860 nm is squeezed by the pump to generate squeezed vacuum (SV). The SV and the remaining pump light enter an adiabatic coupler that passes at least 55% of the SV and 5% of the pump into the measurement OPA. Inside the measurement OPA, the SV is amplified by a strong pump pulse (40 pJ) to macroscopic photon levels detectable by a fast photodetector. A lensed fiber collects light from the output of the measurement OPA and directs it to a series of free-space filters used to remove remaining pump light, including the previously mentioned 1700-nm long pass and a 1950-nm short pass for limiting the measured mode number. From here, the filtered SV is collected into single-mode fiber and enters a 90/10 splitter where 90% is sent to a fast (10 GHz) photodetector and fast (80 GSPS 40 GHz) oscilloscope used to resolve individual pulses while the remaining 10% is sent to a slow (1 MHz) detector and slow (100 MSPS 10 MHz) oscilloscope used to detect the amplification fringe and determine the measurement phase $\phi$.